One side of a rectangle is 14 meters. The perimeter of the rectangle is 44 meters. What is the area of this rectangle?
Answer by Boreal(15235)

half perimeter = L+W
22=14+W
W=8 meters
area is 14 m* 8m =112 m^2
Note the change, 1/2 of the rectangle's perimeter is 22 m
2L+2W=44
L+W=22